Please solve,
Find the length of the sides of a square if the number of square meters in its area is 1.5 timea the number of meters in its perimeter.
Thank you

Answer by ankor@dixie-net.com(12701)

Find the length of the sides of a square if the number of square meters in its area is 1.5 times the number of meters in its perimeter.
:
Let x = side of the square
:
x^2 = 1.5(4x)
x^2 = 6x
x^2 - 6x = 0
x(x-6) = 0
x = 6 meters